# Disentangling CLIP for Multi-Object Perception

### v3

@@ 4.1 Datasets and Metrics
1) Pre-training with MLR: We evaluate the MLR performance using mean-Average Precision (mAP) on datasets: (1) COCO-14 [19] contains 80 classes with 82,081 training and 40,504 validation images. Following recent works [27, 24, 23], we train on the training set and evaluate on the validation set. (2) VOC2007 [11] is another widely used MLR dataset containing 20 classes with 9,963 images. Following [27, 24, 23], we use the train-val set for training and the test set for evaluation.
2) Zero-Shot Semantic Segmentation (ZS3): We use image and text projectors trained on the COCO-14 dataset and evaluate ZS3 using the mIoU metric on the following datasets: (1) PASCAL VOC 2012 [11] includes segmentation masks for the 20 classes in VOC2007. Following works [18, 3], we evaluate on the validation set. (2) COCO-2017 [19] includes segmentation masks for the 80 classes in COCO-14. Following [18, 3], we evaluate on the validation set.
@@ 4.2 Implementation Details
We use CLIP’s [22] original pre-trained encoder weights for all our experiments and keep them frozen. Consistent with popular MLR and ZS3 literature, we use a ResNet-based visual encoder and the standard transformer for text encoding [27, 9, 15, 23, 24, 12, 18, 20]. We conduct all experiments on a single RTX A4000 GPU.
During the pre-training stage with the MLR setup (Sec. 3.4), we follow the settings and hyperparameters from recent works [27, 24, 23]. This includes resizing images to $448$, applying Cutout [8] and RandAugment [7] for augmentation. Our projectors ($h_{\phi}$) are implemented as multi-layer perceptrons (MLPs). Specifically, the image projector follows a [512 $\rightarrow$ 256] architecture, while the text projector is designed as [512 $\rightarrow$ 384 $\rightarrow$ 256] with batch normalization and ReLU. We train both projectors with stochastic gradient descent (SGD) using an initial learning rate of 0.002, which is reduced by cosine annealing. We train the DCLIP setup (ASL + MFI loss) for 50 epochs with a batch size of 32. We follow [27, 24, 23], and use ASL hyperparameters in eq. 2 as $\gamma_{-}=2$, $\gamma_{+}=1$ and $\delta$ = 0.05. We set $\lambda$ = 0.2 and $\alpha$ = $7\mathrm{e}{-5}$ when pre-trained with COCO-14 in eq. 1.
For Zero-Shot Semantic Segmentation, we adopt the v-v attention described in [18] that prevents inversion of activation commonly observed in CLIP. We then add our pre-trained projectors to CLIP. To obtain the segmentation mask, we compute the cosine similarity between locally projected image features ($\mathbf{z^{\prime}}$) and projected text features for all classes in the dataset. We use the template “A photo of a {classname}.” Lastly, we use bilinear interpolation to upsample the mask to the input image size.
Table 2: Comparison on zero-shot semantic segmentation (ZS3). We compare DCLIP with other SOTA baselines across three semantic segmentation datasets using the mIoU metric. The abbreviations are as follows: Loc Ann. + FT: local annotations and fine-tuning, Bkgd: include background class, No Bkgd: ignore background class, MR: MFI Reduction(%), red and blue indicate the best and the second best performance.
@@
SPNet [28] 15.6 - -
ZS3Net [4] 17.7 - -
CLIP-ES [20] 75.0 - -
CLIP [22] 14.1 3.9 5.6
CLIPSurgery [18] 17.5 13.0 22.9
CLIP-VV [18] 32.6 19.9 35.5
Ours (MR = 24.9) 36.0 22.7 37.8
Table 3: Leveraging DCLIP Segments for Enhanced Zero-Shot Multi-Label Recognition. CLIP recognition capabilities often degrade in images with multiple objects. We demonstrate that incorporating object segments generated from DCLIP provides additional evidence to the global image features for zero-shot multi-label recognition (ZS-MLR) on VOC2007 and COCO-14. The resulting performance gains highlight DCLIP’s ability to extract single-object information, enabling robust zero-shot recognition with CLIP in multi-object images.
Dataset Backbone CLIP (mAP) Ours (mAP)
VOC2007 RN 101 78.73 80.71
RN 50 76.20 79.87
COCO-14 RN 101 50.10 52.00
RN 50 47.30 50.15
@@ 4.3 Results
Multi-Label Recognition. We primarily compare DCLIP with other SOTA VLM-based MLR approaches. In Tab. 1, we present a detailed comparison of the performance (mAP-averaged over five runs), number of training parameters, and GPU hours required by each method on the VOC2007 [11] and COCO-14 [19] datasets. For VOC 2007, we observe that DCLIP performs favorably over DualCoOp++[15], requiring equal parameters but fewer training hours. Additionally, on the more challenging COCO-14 dataset, DCLIP outperforms DualCoOp++ while requiring 75% fewer training parameters and fewer training hours on an NVIDIA A4000 GPU.
Zero-Shot Semantic Segmentation. We categorize our comparisons into two groups. The first group includes approaches that use local annotations (segmentation masks) to fine-tune the network [28, 4, 20]. The second group do not use any local annotations [22, 18]. Our approach belongs to the second group, as we do not use any form of local annotations Our results are summarized in Tab. 3. Following prior works [28, 4, 20, 18], we report mIoU values for VOC2012 by including the background as a class. Following [3], we use a threshold of 0.85 to identify the background. Our approach performs favorably over CLIP Surgery by 18.5 mIoU and CLIP-VV by 3.4 mIoU on VOC2012. On COCO-14, our method outperforms CLIP Surgery and CLIP-VV by 9.7 and 2.8 mIoU with background included, and by 14.9 and 2.3 mIoU without background, respectively.
@@ 5 Analysis
Image Oracle CLIP [22] CS [18] CLIP-VV [18] DCLIP (ours)
@@
Figure 5: Qualitative ZS3 Comparison. Visualization of ZS3 results for CLIP [22], CLIP Surgery (CS) [18], CLIP-VV [18], and DCLIP (ours) across multiple categories. The red regions in Oracle show the queried classes. The heatmaps for different methods show activation regions for each queried class, where darker red indicates strongly activated regions. DCLIP (ours) produces more separated activations, leading to improved class localization.
@@
Feature Disentanglement and its Impact. This section examines both quantitative and qualitative impacts of class feature disentanglement. In Fig. 2, we visualize self-similarity matrices for class text features, revealing that DCLIP achieves substantially lower inter-class similarity (off-diagonal values) than baseline CLIP. The distribution of similarity values in Fig. 4 further illustrates how DCLIP shifts feature representations toward reduced inter-class similarity. We quantitatively assess this disentanglement in Tab. 5 using average inter-class similarity, where DCLIP consistently demonstrates lower values across datasets, confirming effective feature separation. Crucially, Fig. 6 establishes a clear inverse relationship between MFI and performance: as MFI decreases, indicating greater feature disentanglement, we observe gains in both multi-label recognition and zero-shot semantic segmentation. Our ablation experiments provide additional validation, showing that removing the MFI Loss component for separation results in a significant 0.6 mAP reduction in MLR task performance (Tab. 5). These results strongly confirm that disentangling class features is both necessary and beneficial for improving performance for multi-object perception.
@@
Segmentation-driven Zero-Shot Multi-Label Recognition (ZS-MLR). To evaluate whether DCLIP produces meaningful segments, we reformulate the multi-label recognition problem into a single-label recognition problem, a domain more suitable for CLIP. Specifically, we use two predictions: global and local. We pass the input image directly through CLIP to obtain its global predictions. However, as discussed in Sec. 1, these predictions are often dominated by more prominent objects in the image, ignoring smaller objects, which leads to poor zero-shot MLR performance. To address this limitation, we introduce local predictions. We segment the image into multiple regions (ideally corresponding to individual objects) using DCLIP. Each segment is then processed independently through CLIP, and the predictions from all segments are aggregated. Finally, we combine the global and local predictions to obtain the zero-shot scores for the image. We evaluate the ZS-MLR performance on the VOC2007 and COCO-14 datasets, with results presented in Tab. 3. The results demonstrate that DCLIP provides meaningful information (individual object segments) to improve CLIP’s zero-shot multi-label recognition capabilities.
@@
Ablation. We study the sensitivity of MLR performance to the sensitivity of hyperparameters and projector choices. (1) $\boldsymbol{\alpha}$ (ASL–MFI Trade-off): Fig. 7 shows that varying $\alpha$, the coefficient that balances the ASL and MFI loss, has minimal impact on performance, indicating low sensitivity to $\alpha$. (2) $\boldsymbol{\lambda}$ (MFI Loss Weighting): We varied $\lambda$, which controls the trade-off between collapse prevention and redundancy reduction in the MFI loss (Eq. 1), across the range $[0.02,0.2]$. Performance remained stable throughout, suggesting insensitivity to $\lambda$ (Appendix Sec E). (3) Projector Dimension: Increasing the dimension of the projector beyond 256 dimensions led to no notable performance gains (Fig. 7).
Table 4: Quantitative MFI Reduction. MFI values are reported across VOC and COCO datasets. DCLIP significantly reduces MFI.
Method VOC COCO
CLIP 0.77 0.69
DCLIP(Ours) 0.50 0.52
$\Delta$ (%) 34.8 24.9
Table 5: MFI Loss Ablation. Without MFI Loss, we observe multi-label recognition performance drops by 0.6 mAP on the COCO-14 dataset.
Method ASL Loss MFI Loss mAP
Ours 84.8
85.4
@@
Figure 6: Performance vs. MFI Reduction. Increase in class feature separation (i.e., MFI decreases) improves performance on multi-label recognition (mAP) and zero-shot semantic segmentation (mIoU) tasks.
@@
Figure 7: Hyperparameter Ablations. 1) DCLIP demonstrates stability across $\alpha$ variations with fluctuations within 2mAP. 2) Performance improves with increased projector dimensionality, saturating beyond 256 dimensions.
6 Conclusions and Broader Impact

### v4

@@ 4.1 Datasets and Metrics
1) Adaptation with MLR: We evaluate the MLR performance using mean-Average Precision (mAP) on datasets: (1) COCO-14 (80 classes) (Lin et al., 2014). Following recent works (Sun et al., 2022; Rawlekar et al., 2024; 2025), we train on the training set and evaluate on the validation set. (2) VOC2007 (20 classes) (Everingham et al., 2010). Following (Sun et al., 2022; Rawlekar et al., 2024; 2025), we use the train-val set for training and the test set for evaluation.
2) Zero-Shot Semantic Segmentation (ZS3): We extract projectors trained for MLR on COCO-14 (image-level labels) and evaluate ZS3 using the mIoU metric on the following datasets: PASCAL VOC 2012 (20 classes + background) (Everingham et al., 2010), COCO-2017(80 classes + background) (Lin et al., 2014), cityscapes (30 classes) (Cordts et al., 2016), context (59 classes) (Mottaghi et al., 2014), stuff (91 classes) (Caesar et al., 2018) and ADE20k (150 classes) (Zhou et al., 2017)
@@ 4.2 Implementation Details
We use CLIP’s (Radford et al., 2021) original pre-trained encoder weights for all our experiments and keep them frozen. Consistent with popular MLR and ZS3 literature, we use a ResNet-based visual encoder and the standard transformer for text encoding (Sun et al., 2022; Ding et al., 2023; Hu et al., 2023; Rawlekar et al., 2025; 2024; Guo et al., 2023; Li et al., 2025; Lin et al., 2023). We conduct all experiments on a single RTX A4000 GPU. For MLR(Sec. 3.4) we follow the settings and hyperparameters from recent works (Sun et al., 2022; Rawlekar et al., 2024; 2025). This includes resizing images to $448$, applying Cutout (DeVries & Taylor, 2017) and RandAugment (Cubuk et al., 2020) transforms. Our projectors ($h_{\phi}$) are implemented as multi-layer perceptrons. Specifically, the image projector follows a [512 $\rightarrow$ 256] architecture, while the text projector is designed as [512 $\rightarrow$ 384 $\rightarrow$ 256] with batch normalization and ReLU. We train both projectors with stochastic gradient descent with an initial learning rate of 0.002, which is reduced by cosine annealing. We train the DCLIP for 50 epochs with a batch size of 32. We follow (Sun et al., 2022; Rawlekar et al., 2024; 2025), and use ASL hyperparameters in eq. 5 as $\gamma_{-}=2$, $\gamma_{+}=1$ and $\delta$ = 0.05. We set $\lambda$ = 0.2 and $\alpha$ = $7\mathrm{e}{-5}$ when pre-trained with COCO-14 in eq. 2. For Zero-Shot Semantic Segmentation, we adopt the vv attention (Li et al., 2025) that prevents inversion of activation commonly observed in CLIP. We then add our pre-trained projectors to CLIP. To obtain the segmentation mask, we compute the cosine similarity between locally projected image features ($\mathbf{z^{\prime}}$) and projected text features for all classes in the dataset. We use the template “A photo of a {classname}.” Lastly, we use bilinear interpolation to upsample the mask to the input image size.
Table 2: Zero-shot semantic segmentation (ZS3) comparison. We compare DCLIP with other SOTA methods using mIoU metric. The abbreviations are: Loc Ann.+ FT: local annotations and fine-tuning, Bkgd: include background class, No Bkgd: ignore background, red and blue indicate the best and the second best performance.
@@
SPNet (Xian et al., 2019) 15.6 - -
ZS3Net (Bucher et al., 2019) 17.7 - -
CLIP-ES (Lin et al., 2023) 75.0 - -
CLIP (Radford et al., 2021) 14.1 3.9 5.6
CLIPSurgery (Li et al., 2025) 17.5 13.0 22.9
CLIP-VV (Li et al., 2025) 32.6 19.9 35.5
36.0 22.7 37.8
Table 3: MFI Reduction. MFI values for VOC and COCO. DCLIP significantly reduces MFI.
Method VOC COCO
CLIP 0.77 0.69
DCLIP 0.50 0.52
$\Delta$ (%) 34.8 24.9
Table 4: MFI Loss Ablation. Without MFI, MLR performance drops by 1.2 mAP on COCO.
Method ASL MFI mAP
Ours 84.2
85.4
@@ 4.3 Results
Multi-Label Recognition. We compare DCLIP with other SOTA VLM-based MLR approaches. In Tab. 1, we present a detailed comparison of the performance (mAP-averaged over five runs), number of training parameters, and GPU hours required by each method on the VOC2007 (Everingham et al., 2010) and COCO-14 (Lin et al., 2014) datasets. For VOC 2007, we observe that DCLIP performs favorably over DualCoOp++(Hu et al., 2023), requiring equal parameters but fewer training hours. Additionally, on the more challenging COCO-14 dataset, DCLIP outperforms DualCoOp++ while requiring 75% fewer training parameters and fewer training hours on an NVIDIA A4000 GPU.
Zero-Shot Semantic Segmentation. We categorize our comparisons into two groups. The first group includes approaches that use local annotations (segmentation masks) to fine-tune the network (Xian et al., 2019; Bucher et al., 2019; Lin et al., 2023). The second group does not use any local annotations (Radford et al., 2021; Li et al., 2025). As we do not use any form of local annotations, DCLIP belongs to the second group. Our projectors train only on image-level MLR labels. Results are summarized in Tab. 4.2. For VOC2012, we report the mIoU with background. Following (Bousselham et al., 2024), we use a threshold of 0.85 to identify the background. Our approach performs favorably over CLIP Surgery by 18.5 mIoU and CLIP-VV by 3.4 mIoU on VOC2012. On COCO-17, our method outperforms CLIP Surgery and CLIP-VV by 9.7 and 2.8 mIoU with background, and by 14.9 and 2.3 mIoU without background. Fig. 6 shows performance on cityscapes (Cordts et al., 2016), context (Mottaghi et al., 2014), stuff (Caesar et al., 2018) and ADE20k (Zhou et al., 2017). This demonstrates DCLIP’s ability to learn domain-agnostic disentangled features that transfer across diverse visual environments (urban scenes, indoor scenes, natural images).
@@ 5 Analysis
Image Oracle CLIP Radford et al. (2021) CS Li et al. (2025) CLIP-$V$ Li et al. (2025) DCLIP (ours)
@@
Figure 5: Qualitative ZS3 Comparison. Visualization of ZS3 results for CLIP (Radford et al., 2021), CLIP Surgery (CS) (Li et al., 2025), CLIP-VV (Li et al., 2025), and DCLIP (ours) across multiple categories. The red regions in Oracle show the queried classes. The heatmaps for different methods show activation regions for each queried class, where darker red indicates strongly activated regions. DCLIP (ours) produces more separated activations, leading to improved class localization.
@@
Feature Disentanglement and its Impact. This section examines both quantitative and qualitative impacts of class feature disentanglement. In Fig. 2, we visualize self-similarity matrices for class text features, revealing that DCLIP achieves substantially lower inter-class similarity (off-diagonal values) than baseline CLIP. The distribution of similarity values in Fig. 4 further illustrates how DCLIP shifts feature representations toward reduced inter-class similarity. We quantitatively assess this disentanglement in Tab. 4.2 using average inter-class similarity, where DCLIP consistently demonstrates lower values across datasets, confirming effective feature separation. Crucially, Fig. 7 establishes a clear inverse relationship between MFI and performance: as MFI decreases, indicating greater feature disentanglement, we observe gains in both multi-label recognition and zero-shot semantic segmentation. Our ablation experiments provide additional validation, showing that removing the MFI Loss component for separation results in a significant 1.2 mAP reduction in MLR task performance (Tab. 4.2). These results strongly confirm that disentangling class features is both necessary and beneficial for improving performance for multi-object perception.
@@
DCLIP’s Segments for Multi-Label Recognition. To evaluate whether DCLIP produces meaningful object segments, we test its utility in improving CLIP’s MLR performance. We reformulate the MLR problem by combining global and local predictions. For global predictions, we pass the input image directly through the original CLIP model. However, as discussed in Sec. 1, these predictions are often dominated by more prominent objects, ignoring smaller objects in multi-object scenes. To complement this, we generate local predictions by first segmenting the image using DCLIP to isolate individual objects, then processing each segment independently through the original CLIP model. The predictions from all segments and global image are combined to get final scores. The results in Tab. 5 on VOC2007 and COCO-14 demonstrate that DCLIP extracts meaningful object segments.
@@
Ablation. We study the sensitivity of MLR performance to the sensitivity of hyperparameters and projector choices. (1) $\bm{\alpha}$ (ASL–MFI Trade-off): Fig. 8 shows that varying $\alpha$, the coefficient that balances the ASL and MFI loss, has minimal impact on performance, indicating low sensitivity to $\alpha$. (2) $\bm{\lambda}$ (MFI Loss Weighting): We varied $\lambda$, which controls the trade-off between collapse prevention and redundancy reduction in the MFI loss (Eq. 2), across the range $[0.02,0.2]$. Performance remained stable throughout, suggesting insensitivity to $\lambda$ (Appendix Sec D.1). (3) Projector Dim: Increasing the dimension of the projector beyond 256 dim led to saturation (Fig. 8). Ablations on loss (BCE vs Focal vs ASL), pooling–projection order and architectures are in Appendix Sec. D.2, D.3, D.4.
Table 5: CLIP’s Multi-Label Recognition with DCLIP Segments. By segmenting images into individual objects using DCLIP and processing each segment through CLIP, we achieve consistent improvements on VOC2007 and COCO-14.
Dataset Backbone CLIP (mAP) + DCLIP Segments (mAP) VOC2007 RN101 78.73 80.71 RN50 76.20 79.87 COCO-14 RN101 50.10 52.00 RN50 47.30 50.15
Figure 6: ZS3 results for Cityscapes, Pascal Context, Stuff and ADE20K dataset.
@@
Figure 7: Performance vs. MFI Reduction. Increase in class feature separation (i.e., MFI decreases) improves performance on MLR (mAP) and ZS3 (mIoU) tasks.
@@
Figure 8: Hyperparameter Ablations. 1) DCLIP demonstrates stability across $\alpha$ variations. 2) Increasing projector dimension improves performance (saturates beyond 256 dims).
6 Conclusions